lance is now twice as old as gean. six years ago, the sum of their ages was 9. how old is lance now?
Answer by TimothyLamb(4379)

x = lance age now
y = gean age now
---
x = 2y
x - 6 + y - 6 = 9
---
put the system of linear equations into standard form
---
x - 2y = 0
x + y = 21
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = lance age now = 14
y = gean age now = 7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php